# Finite Dynamical Laminations

### v2

@@ 4 Finite Dynamic Laminations
A lamination $\mathcal{L}$ is called a Finite Dynamical Lamination or an FDL, in degree $d$ if it is non-empty and satisfying these properties:
@@
Given that the endpoints of all the leaves of such a lamination are eventually periodic, those endpoints are rational angles. Since [3] establishes that all rational rays land the following definition is unambiguous. Moreover, we will show that it is a non-empty set (Theorem 6.3).
@@
The preceding sections first introduce the pullback tree, and then later the generational FDL graph. The latter is perhaps the more illuminating structure for understanding parameter space, but its vertices are those from a single level of the former.
It is possible to surjectively parameterize the set of FDL, with a natural number, $n$, a set of periodic leaves, $P$, and a pullback scheme. We take a slice of the set of FDL, by specifying $P$.
@@
Figure 3: The pullback tree starting with the polygon with vertices {_$\_001$,_$\_010$,_$\_100$} up to level 5, also known as the rabbit pullback tree.
@@ 5 What are the limit sets of FDL?
The goal of this section is to establish some facts about the limit of a sequence of FDL. Nothing good follows from the definition of FDL. The set of finite dynamical laminations is not closed and contains none of its limits. The term “limit lamination” has an existing meaning and therefore [9], the limit of a sequence of FDL will be referred to as an “end lamination”. First, we situate FDL inside a larger set of laminations. Then after considering how bad sequences of FDL are in general, we impose a fairly obvious restriction, and find that such sequences limit to reasonably nice laminations.
Of course, we wish to classify these sequences by how nice their limits are. For that purpose we summarize the definitions of a few terms which we use sparsely. The most important term is q-lamination, which is defined in [1] as a lamination having an equivalent relation that is laminational with the added convention that a leaf is in the lamination iff it is in the boundary of the convex hull of an equivalence class. Laminational, is also defined in [1], and refers to an equivalence relation that meets a number of conditions that are each necessary for the lamination to be realized by a polynomial in the sense contemplated here. One such condition is that the equivalence classes are finite. These terms generally presuppose that the lamination is invariant. Unclean is defined in Thurston as having 3 leaves meet at a point. Unclean laminations are not q, but they might be proper, in which case their equivalence relations are laminational. When we say that a lamination is not proper we mean that it has infinite equivalence classes. To summarize, an unclean sibling lamination can be cleaned into a q-lamination iff it is proper.
In our examples, which are quadratic, we give a few more specific labels mainly from [9]. We say that a lamination is hyperbolic if there is a critical gap with uncountably many sides. If there is an uncountable gap that returns with degree 1, then we say that it and the lamination are Siegel. We call a gap caterpillar if it has countably many leaves. In accordance with folklore, we say that a lamination is sub-hyperbolic if there is a finite equivalence class that is critical and preperiodic.
@@
The maximum distance from one lamination to another is at most the greatest distance from one leaf to another which is $\frac{1}{2}$. As for the closed part of the claim, suppose by contradiction that the limit of a sequence of laminations had a pair of crossing leaves. The arcs subtended by those leaves overlap in an arc of length $\epsilon>0$. Select from the sequence, a lamination such that every leaf of the limit has a leaf of the lamination within $\frac{\epsilon}{3}$. The leaves of the lamination close to the crossing leaves in the limit would have to cross because the arcs subtending them must have an intersection of length at least $\frac{\epsilon}{3}$. ∎
@@
Of course a sequence of FDL might be a constant sequence or contain a constant subsequence. Suppose it does not. It might oscillate. Suppose it converges to a lamination not in the sequence. Surely then it will be an invariant lamination?
@@
There is a non-constant converging sequence of FDL whose limit is not invariant. Consider the canonical rabbit lamination, and consider the triangles on the boundary of the image of the critical gap. Form a converging sequence of those triangles. Preferably the limit of this sequence should be the singleton at _$\_010$. Next, for each triangle in the sequence, form the minimal FDL containing the hexagon that maps 2 to 1 onto that triangle. The limit of this sequence of FDL is the rabbit lamination with a critical leaf touching the periodic point. But that critical leaf has no preimage.
@@
Figures 7 and 8 are the results of forcing the lamination to place a leaf at a certain point in the manner contemplated above. In Figure 7, we force it to place a limit leaf at $0010\_001$. In the FDL of the sequence, there are already 2 other leaves there for a total of 3, making the lamination not clean. However, the lamination is not too bad and amounts to another way to approach the Misiurewicz point.
Figure 8 is the result of forcing the lamination to have a leaf with the endpoint _$\_0010$. If you have very good eyes you may be able to discern that the critical gap is a Fatou gap. This is an example of achieving a primitive bifurcation and forcing a periodic leaf to appear. It seems that at many points we can force there to be a leaf, but the lamination chooses what kind of leaf.
@@ 6 Are all FDL realized by a polynomial?
Though the next logical step may be to contemplate loops in the closure of the pullback tree, a patient reader may be interested in whether all FDL have polynomials, and we are ready to answer that. After we take a moment to observe the absence of Siegel gaps, we can use Kiwi’s theorem [13]. Kiwi defined the term “$\mathbb{R}$eal lamination” to refer to laminational equivalence relations. [1] states that forward invariance implies backward invariance. Otherwise, the stipulations in the definitions are in one to one correspondence with identical meaning. Thus, we use the term laminational as in [1].
@@
Kiwi’s theorem does not show that all rays land in the way described by the lamination but instead that the impression of each ray intersects as described by the lamination. Along with the fact that all FDL have at least one child, this is enough to establish that the co-landing locus of any FDL is non-empty. The vertices of polygons in an FDL are all periodic or preperiodic, which is the same as rational. According to [3], all rational rays land at periodic or preperiodic point in the Julia set. Kiwi makes it clear with his characterization of what it means for a ray to land that the landing point of a ray is in its impression. Thus, his corollary 1.2 applies, and the impression of the ray is its landing point. ∎
@@ 7 Loops in the closure of the pullback tree?
By Lemma 2.13, the free criticality and the trapped criticality add to $d-1$.
@@ 8 Future Directions of Research
    Consider the quadratic pullback tree starting with the lamination containing only the leaf from _$\_01$ to _$\_10$. The number of FDL on each level of the tree is A152046 in [8].

### v3

@@ 4 Finite Dynamic Laminations
A lamination $\mathcal{L}$ is called a Finite Dynamical Lamination or an FDL, in degree $d$ if it is non-empty and satisfies these properties:
@@
Given that the endpoints of all the leaves of such a lamination are eventually periodic, those endpoints are rational angles. Regarding the following definition, land is taken to have the same meaning as in [3], which establishes that a rational ray will always either land or crash into a critical point. We will show that the set defined is a non-empty set (Theorem 6.3).
@@
The following sections first introduce the pullback tree, and then later the generational FDL graph. The latter is perhaps the more illuminating structure for understanding parameter space, but its vertices are those from a single level of the former.
It is possible to surjectively parameterize the set of FDL, with a natural number, $n$, a set of periodic leaves, $P$, and a pullback scheme. We take a slice of the set of FDL by specifying $P$.
@@
Figure 3: The pullback tree starting with the polygon with vertices {_001\_001,_010\_010,_100\_100} up to level 5, also known as the rabbit pullback tree.
@@ 5 What are the limit sets of FDL?
The goal of this section is to establish some facts about the limit of a sequence of FDL. Nothing good follows from the definition of FDL. The set of finite dynamical laminations is not closed and contains none of its limits. The term “limit lamination” has an existing meaning[8], so the limit of a sequence of FDL will be referred to as an “end lamination.” First, we situate FDL inside a larger set of laminations. Then after considering how bad sequences of FDL are in general, we impose a fairly obvious restriction and find that such sequences limit to reasonably nice laminations.
Of course, we wish to classify these sequences by how nice their limits are. For that purpose, we summarize the definitions of a few terms that we use sparsely. The most important term is q-lamination, which is defined in [1] as a lamination having an equivalent relation that is laminational with the added convention that a leaf is in the lamination iff it is in the boundary of the convex hull of an equivalence class. Laminational is also defined in [1] and refers to an equivalence relation that meets several conditions that are each necessary for the lamination to be realized by a polynomial in the sense contemplated here. One such condition is that the equivalence classes are finite. These terms generally presuppose that the lamination is invariant. Unclean is defined in Thurston as having 3 leaves meet at a point. Unclean laminations are not q-laminations, but they might be proper, in which case their equivalence relations are laminational. When we say that a lamination is not proper, we mean that it has infinite equivalence classes. To summarize, an unclean sibling lamination can be cleaned into a q-lamination iff it is proper.
In our examples, which are quadratic, we give a few more specific labels, mainly from [8]. We say that a lamination is hyperbolic if there is a critical gap with uncountably many sides. If there is an uncountable gap that returns with degree 1, then we say that it and the lamination are Siegel. We call a gap caterpillar if it has countably many leaves. In accordance with folklore, we say that a lamination is sub-hyperbolic if there is a finite equivalence class that is critical and preperiodic.
@@
The maximum distance from one lamination to another is at most the greatest distance from one leaf to another, which is $\frac{1}{2}$. As for the closed part of the claim, suppose by contradiction that the limit of a sequence of laminations had a pair of crossing leaves. The arcs subtended by those leaves overlap in an arc of length $\epsilon>0$. Select from the sequence, a lamination such that every leaf of the limit has a leaf of the lamination within $\frac{\epsilon}{3}$. The leaves of the lamination close to the crossing leaves in the limit would have to cross because the arcs subtending them must have an intersection of length at least $\frac{\epsilon}{3}$. ∎
@@
Of course, a sequence of FDL might be a constant sequence or contain a constant subsequence. Suppose it does not. It might oscillate. Suppose it converges to a lamination not in the sequence. Surely then it will be an invariant lamination?
@@
There is a non-constant converging sequence of FDL whose limit is not invariant. Consider the canonical rabbit lamination, and consider the triangles on the boundary of the image of the critical gap. Form a converging sequence of those triangles. Preferably the limit of this sequence should be the singleton at _010\_010. Next, for each triangle in the sequence, form the minimal FDL containing the hexagon that maps 2 to 1 onto that triangle. The limit of this sequence of FDL is the rabbit lamination with a critical leaf touching the periodic point. But that critical leaf has no preimage.
@@
Figures 7 and 8 are the results of forcing the lamination to place a leaf at a certain point in the manner contemplated above. In Figure 7, we force it to place a limit leaf at 0010_0010010\_001. In the FDL of the sequence, there are already 2 other leaves there for a total of 3, making the lamination not clean. However, the lamination is not too bad and amounts to another way to approach the Misiurewicz point.
Figure 8 is the result of forcing the lamination to have a leaf with the endpoint _0010\_0010. If you have very good eyes you may be able to discern that the critical gap is a Fatou gap. This is an example of achieving a primitive bifurcation and forcing a periodic leaf to appear. It seems that at many points we can force there to be a leaf, but the lamination chooses what kind of leaf.
@@ 6 Are all FDL realized by a polynomial?
Though the next logical step may be to contemplate loops in the closure of the pullback tree, a patient reader may be interested in whether all FDL have polynomials, and we are ready to answer that. After we take a moment to observe the absence of Siegel gaps, we can use Kiwi’s theorem [12]. Kiwi defined the term $\mathbb{R}$eal lamination” to refer to laminational equivalence relations. [1] states that forward invariance implies backward invariance. Otherwise, the stipulations in the definitions are in one to one correspondence with identical meaning. Thus, we use the term laminational as in [1].
@@
Kiwi’s theorem does not show that all rays land in the way described by the lamination but instead that the impression of each ray intersects as described by the lamination. Along with the fact that all FDL have at least one child, this is enough to establish that the co-landing locus of any FDL is non-empty. The vertices of polygons in an FDL are all periodic or preperiodic, which is the same as rational. Using the connectedness of the Julia set, we can see from [3] that the rational rays land. Also from [3], we know all rational rays land at periodic or preperiodic point in the Julia set. Kiwi makes it clear with his characterization of what it means for a ray to land that the landing point of a ray is in its impression. Thus, his corollary 1.2 applies, and the impression of the ray is its landing point. ∎
@@ 7 Loops in the closure of the pullback tree?
By Lemma 2.13, the free criticality and the trapped criticality add to d1d-1.
@@ 8 Future Directions of Research
    Consider the quadratic pullback tree starting with the lamination containing only the leaf from _01\_01 to _10\_10. The number of FDL on each level of the tree is A152046 in [7].
@@
    Regarding the closed part, Proposition 5.1 [14] states that the landing point of a rational ray of a quadratic polynomial will very continuously with the parameter. Hopefully, this can be generalized.